OpenStudy (lgbasallote):

i don't know if this is a hard one...but it LOOKS complicated ;D Prove that: \[\mathbf{\frac {cos (7A) + cos (3A) - cos (5A) - cos(A)}{sin(7A) - sin(3A) - sin (5A) + sin(A)} = cot(2A)}\]

OpenStudy (lgbasallote):

is it o.O

OpenStudy (diyadiya):

Identity

OpenStudy (mimi_x3):

It's not a formula..you have to like prove it..there's no formula for cot(x+y) i guess.

OpenStudy (lgbasallote):

nahh i'll just change it...if it's like a formula then it's easy

OpenStudy (diyadiya):

Its an IDENTITY!!

OpenStudy (blockcolder):

Maybe you derive it the same way you derive tan(x+y).

OpenStudy (wasiqss):

yeh we need to open by the formula {(cos(x+y)}/sin(x+y)}

OpenStudy (diyadiya):

Yeah ^

OpenStudy (lgbasallote):

i think that's not a formula anymore

OpenStudy (diyadiya):

LOl you can't change the question like that

OpenStudy (lgbasallote):

fine i'll post in a different question :P

OpenStudy (diyadiya):

No need

OpenStudy (mimi_x3):

lol, that is a very long proof..will take ages to type..

OpenStudy (diyadiya):

Using cosx+cosy & cosx-cosy identity

OpenStudy (diyadiya):

then sinx-siny & sinx+siny

OpenStudy (lgbasallote):

long proof..ha! i knew it was hard \m/

OpenStudy (mimi_x3):

probably double angle formulae..

OpenStudy (wasiqss):

its not hard, its just we dont have such time to waste on the latex :P

OpenStudy (mimi_x3):

like wasoo knows how to use latex. :P

OpenStudy (wasiqss):

hehe thats another case :P but i learned it :P

OpenStudy (mimi_x3):

lol, then use it to solve this problem. :P

OpenStudy (wasiqss):

across taught me i guess :P

OpenStudy (diyadiya):

sinX + sinY = 2sin[ (X + Y) / 2 ] cos[ (X - Y) / 2 ] sinX - sinY = 2cos[ (X + Y) / 2 ] sin[ (X - Y) / 2 ] cosX + cosY = 2cos[ (X + Y) / 2 ] cos[ (X - Y) / 2 ] cosX - cosY = - 2sin[ (X + Y) / 2 ] sin[ (X - Y) / 2 ]

OpenStudy (lgbasallote):

\[\huge \mathbf{L} \mathsf{A} \mathbf{T} \mathsf{E} \mathbf{X?}\]

OpenStudy (lgbasallote):

hehe =))

OpenStudy (lgbasallote):

@Diyadiya are you taking trig now? for you to remember those...

OpenStudy (diyadiya):

i've been doing trig since last 3yrs :D

OpenStudy (mimi_x3):

\[\cot2A = \frac{\cos2A}{\sin2A} =\frac{2\cos^{2}A-1}{2sinAcosA} \] I'm doubting that it is provable..

OpenStudy (diyadiya):

IDK .. LGB Type down what you did till now :D

OpenStudy (lgbasallote):

proving is hard isnt it @_@

OpenStudy (lgbasallote):

i mean tiresome

OpenStudy (mimi_x3):

Looks like it would work..but will take a while..

OpenStudy (mimi_x3):

any ideas diya?

OpenStudy (mimi_x3):

well, we know that: \(\cos3x = 4\cos^{2} x-3cosx\) but what about cos(7x)?